# Data-driven Model Predictive Control: Asymptotic Stability despite Approximation Errors exemplified in the Koopman framework

### v1

@@ 4.2. Extension of kernel EDMD to control systems
$c_{x}^{\varepsilon}=c_{x}^{h}h_{\mathcal{X}}^{k-1/2}+c_{x}^{r}r_{\mathcal{X}},\qquad c_{u}^{\varepsilon}=c_{u}^{h}h_{\mathcal{X}}^{k+1/2}+c_{u}^{r}r_{\mathcal{X}}$
for suitable $c_{x}^{h},c_{x}^{r},c_{u}^{h},c_{u}^{r}$. The terms $c_{x}^{r}$ and $c_{u}^{r}$ have a dependence on $\|K_{\mathcal{X}}^{-1}\|$, which is related to the location and number $d$ of virtual observation points. For increasing $d$, the minimal eigenvalue of the kernel matrix $K_{\mathcal{X}}$ may decrease, and consequently $\|K_{\mathcal{X}}^{-1}\|$ may increase. Hence, to satisfy given bounds for $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$, one should first select the location of the virtual observation points (that is related to the fill distance $h_{\mathcal{X}}$) to obtain sufficiently small terms $c_{x}^{h}h_{\mathcal{X}}^{k-1/2}$ and $c_{u}^{h}h_{\mathcal{X}}^{k+1/2}$. Given the location of the virtual observation points, $\|K_{\mathcal{X}}^{-1}\|$ is fixed. Then, in a second step, it is possible to select the cluster radius $r_{\mathcal{X}}$ sufficiently small to satisfy the required bounds for the terms $c_{x}^{r}r_{\mathcal{X}}$ and $c_{u}^{r}r_{\mathcal{X}}$.
The proportional error bound derived in Proposition 2 may be more conservative than the bounds of Theorem 3 further away from the origin, but converges to zero for $(x,u)\rightarrow(0,0)$, which is a key property to rigorously show asymptotic stability of the MPC, as shown in Theorem 1 and summarized in the following corollary.
Corollary 2.
Let Assumption 5 on the data hold and consider the data-driven MPC scheme summarized in Algorithm 1 using the kEDMD surrogate model (25) of the control-affine dynamics (22) in the prediction and optimization step. Further, assume that the system is cost controllable, i.e. Assumption 4 holds with $\mathbf{u}\in\mathcal{U}_{N}^{\varepsilon}(\hat{x})\subset\mathcal{U}_{N}(\hat{x})$. Then, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ are sufficiently small, Algorithm 1 asymptotically stabilizes the system (22).
Proof. We have verified Assumptions 1, 2 and 3 for the kEDMD surrogate of the control-affine dynamics (22) in Proposition 2, Theorem 3, and Lemma 1, respectively. Hence, all assumptions of Theorem 1 hold in view of the imposed cost controllability and the assertion can be inferred. $\square$
@@ 4.3. Relation to kernel approximation of control systems
$\widetilde{F}_{i}(x)=\sum_{j=1}^{d}(K_{\mathcal{X}}^{-1}(F_{i})_{\mathcal{X}})_{j}\Phi_{x_{j}}(x)=(F_{i})_{\mathcal{X}}^{\top}K_{\mathcal{X}}^{-1}\begin{bmatrix}\Phi_{x_{1}}(x)\\ \vdots\\ \Phi_{x_{d}}(x)\end{bmatrix}$ (30)
for each $i\in[1:n]$. This corresponds to direct interpolation of the flow map and hence involves only one projection. Moreover, the main difference to the approximation (20) are the data requirements: For the approximation (20) applied to the coordinate functions, we require data through the lens of the features, i.e., $\mathsf{k}(x_{i},F(x_{j}))$ for $(i,j)\in[1:d]^{2}$, whereas for the approximation (30), we require samples of the flow map, i.e., $F(x_{j})$ for all $j\in[1:d]$. The error bound for the surrogate dynamics (30), however, is very closely related to the one inferred from (20) and of similar structure, see (kohne2024infty, , Theorem 3.4), where the first estimate in this reference corresponds to (30) (involving only one projection) and the second estimate to (20) (involving two projections). As both projection errors are proportional to the fill distance $h_{\mathcal{X}}$, the convergence rate is same and the results of Subsection 4.2 can be derived analogously. Hence, Assumptions 1 and 2 can also be verified for the surrogate model (30).
@@ 5. Numerical simulations
In this section, we illustrate the asymptotic stability of the origin w.r.t. the MPC closed loop using the kernel-EDMD surrogate in the prediction and optimization step by conducting numerical simulations. In the following, we denote by physics-informed kEDMD (PI-kEDMD) the models obtained with the modified Step 1 for the encoding of the equilibrium as described in Section 4.2, and with kEDMD the models obtained according to BoldPhil24 without modifying Step 1.
@@ 5.1. Van der Pol
with parameters $\Delta t=0.05$, $\nu=0.1$, and control constraints $\mathbb{U}=[-2,2]$ on the domain $S=[-2,2]^{2}$. The kEDMD surrogate (25) is constructed using Wendland kernels with smoothness degree $k=1$, see Remark 1 for details. The cluster points are chosen according to a Padua grid. The set of Padua points of order $p$ in the domain $[-1,1]\times[-1,1]$ is defined as
@@
Moreover, we add a cluster point in the origin to the Padua grid, according to Assumption 5. The Padua grid is properly scaled to cover the domain $S$. Padua grids of order $p\in\{25,50\}$ are considered, resulting in $d\in\{352,1327\}$ state-space sample points. In comparison to a uniform grid, the use of a Padua grid allows to have a more uniform error through the domain, and lower condition numbers in the model matrices. In fact, Padua points minimize, analogously to Chebychev nodes in one dimension, the Lebesgue constant and, thus, the condition number and are, thus, preferable. The use of a uniform grid with a large number of data points may lead to numerical problems in the solution of the optimal control problem.
For each cluster point $x_{i}\in\mathcal{X}$, $d_{i}=25$ random control values $u_{ij}\in\mathbb{U}$ are chosen, yielding the data points $(x_{ij},u_{ij},x_{ij}^{+})$. Herein, $x_{ij}$ and $u_{ij}$ are drawn according to our data requirements with $r_{\mathcal{X}}=\nicefrac{{\sqrt{2}}}{{d}}$ to specify the neighborhood of $x_{i}$, see Section 4. For the derivation of the PI-kEDMD model Step 1 of the algorithm is modified for the cluster point $x_{1}=0$ by setting $\tilde{g}_{0}(x_{1})=0$, while in the kEDMD model Step 1 is performed in the same way for all cluster points. The MPC cost function is chosen as in (6) with matrices $Q=I_{2}$ and $R=10^{-4}$. The following closed-loop simulations emanate from the initial state $x^{0}=(0.5,0.5)^{\top}$, and consider a prediction horizon $N\in\{10,30\}$ in MPC. The comparison of the closed loop error with different number of clusters and different prediction horizons is reported in Fig. 1.
@@
In the simulations with the kEDMD model, we observe practical asymptotic stability only: the error stagnates at a positive constant value. The tracking accuracy increases with the number of cluster points. This is expected, as a large number of cluster points is related to lower modeling errors. On the other hand, the PI-kEDMD models allow to reach a lower error, with a decrease up to the solver tolerance at around $10^{-14}$, in accordance to the shown asymptotic stability property of the MPC closed loop. We also notice that a fast decrease to the origin is obtained with the prolonged prediction horizon $N=30$ – a property related to the degree of suboptimality $\alpha_{N}$. Finally, in Fig. 2, we report the trajectories of the same simulations in the phase space. We can see that, while the asymptotic behavior is strongly related to the accuracy of the model in the origin, the transient behavior is mainly influenced by the prediction horizon and the number of cluster points.
@@ 5.2. Four-tanks process
As a second example, we consider the four-tank system (see alvarado2011comparative ) described by the differential equation
@@
The state of the system is given by the levels in the four tanks, i.e. $x=(h_{1},\;h_{2},\;h_{3},\;h_{4})^{\top}\in\mathbb{R}^{4}$, while the control variables are the flows through the two valves, i.e. $u=(q_{a},\;q_{b})^{\top}\in\mathbb{R}^{2}$. The numerical values of the system parameters can be found in alvarado2011comparative and as sampling time we choose $\Delta t=10s$. The discrete-time version of (32) obtained with the forward Euler method is considered as ground truth. The controlled state for this system is the equilibrium given by $\bar{x}=(0.65,0.66,0.6417,0.6882)^{\top}m$ and $\bar{u}=(1.666,1.974)^{\top}m^{3}/h$. To derive the kEDMD surrogate model, the state and inputs are shifted around the origin. The state domain, where we sample, is $S=[0.2,1.36]^{2}\times[0.2,1.30]^{2}m$, while the input is constrained by the set $\mathbb{U}:=[0,3.26]\times[0,4]m^{3}/h$. The cluster centers are chosen in a uniform grid in $S$. Moreover, an additional cluster center is added at the equilibrium. We consider datasets with $d\in\{626,1297,2402\}$ cluster centers, that correspond to $4]{d-1}\in\{5,6,7\}$. We decided to stick to a uniform grid, because the explicit construction of a grid minimizing the Lebesgue constant in dimensions higher than two is more involved and we did not run into numerical issues. Possible approaches to reduce the condition number of the model matrices are the use of the cartesian product of Padua grids or of more sophisticated approaches as outlined, e.g., in jimenez2025approximating . The cluster radius is $r_{\mathcal{X}}=2/d$ for all the models, and 25 data points $(x_{ij},u_{ij},x_{ij}^{+})$ are are considered in each cluster. For the closed-loop simulations, we consider an initial state $x_{0}=(1.0,1.0,1.0,1.0)^{\top}$, MPC cost matrices $Q=I_{4}$ and $R=10^{-4}I_{2}$, and a prediction horizon $N=10$.
@@
The errors of the closed loop simulations are reported in Fig. 3. As in the previous example, we can see that the error reaches a higher accuracy with the PI-kEDMD model, and the accuracy increases with a larger number of cluster points. The number of cluster points is also related to the convergence speed, that is lower for $d=626$. In this example, the closed loop behavior remains the same if the prediction horizon is increased to $N=20$.
@@ 6. Conclusions and outlook
In this paper, we have studied MPC schemes without terminal conditions, in which data-driven surrogate models are used in the prediction and optimization step. We have proposed a framework to rigorously show asymptotic stability of an equilibrium w.r.t. the MPC closed loop assuming that the original system is cost controllable. The key ingredients are three properties of the data-driven surrogate: First, a uniform Lipschitz continuity w.r.t. the approximation accuracy (see Assumption 3). Second and third, uniform and foremost proportional finite-data error bounds (see Assumptions 1 and 2). Then, cost controllability is preserved and the desired asymptotic stability can be inferred despite the use of the data-driven surrogate for the prediction and optimization in MPC.
Then, we have rigorously verified all required assumptions on data-driven surrogate models generated with kernel EDMD. In particular, we have proposed a learning algorithm based on flexible data samples, that allows to obtain proportional error bounds by leveraging knowledge of the desired (controlled) equilibrium. Finally, the theoretical results are verified by numerical simulations for two representative examples.
Future work could consider an extension towards Koopman-based control of partial differential equations Deut24 as well as an extension towards novel (E)DMD variants such as eigenfunction-based approximation via (control) Liouville operators RoseKama24 , residual DMD ColbAyto23 or via stochastic Koopman approximants WannMezi22 .

### v2

@@ 4.2. Extension of kernel EDMD to control systems
$\begin{array}[]{rll}c_{x}^{\varepsilon}&=c_{x}^{h}h_{\mathcal{X}}^{k-1/2}&+c_{x}^{r}r_{\mathcal{X}}\\ c_{u}^{\varepsilon}&=c_{u}^{h}h_{\mathcal{X}}^{k-1/2}\operatorname{dist(x,\mathcal{X})}&+c_{u}^{r}r_{\mathcal{X}}\end{array}$
for suitable $c_{x}^{h},c_{x}^{r},c_{u}^{h},c_{u}^{r}$. The terms $c_{x}^{r}$ and $c_{u}^{r}$ have a dependence on $\|K_{\mathcal{X}}^{-1}\|$, which is related to the location and number $d$ of virtual observation points. For increasing $d$, the minimal eigenvalue of the kernel matrix $K_{\mathcal{X}}$ may decrease, and consequently $\|K_{\mathcal{X}}^{-1}\|$ may increase. Hence, to satisfy given bounds for $c_{x}^{\varepsilon}$ and $c_{u}^{\varepsilon}$, one should first select the location of the virtual observation points (that is related to the fill distance $h_{\mathcal{X}}$) to obtain sufficiently small terms $c_{x}^{h}h_{\mathcal{X}}^{k-1/2}$ and $c_{u}^{h}h_{\mathcal{X}}^{k-1/2}\operatorname{dist(x,\mathcal{X})}$. Given the location of the virtual observation points, $\|K_{\mathcal{X}}^{-1}\|$ is fixed. Then, in a second step, it is possible to select the cluster radius $r_{\mathcal{X}}$ sufficiently small to satisfy the required bounds for the terms $c_{x}^{r}r_{\mathcal{X}}$ and $c_{u}^{r}r_{\mathcal{X}}$.
The proportional error bound derived in Proposition 2 may be more conservative than the bounds of Theorem 3 further away from the origin, but converges to zero for $(x,u)\rightarrow(0,0)$, which is a key property to rigorously show asymptotic stability of the MPC, as shown in Theorem 1 and summarized in the following theorem.
Theorem 4.
Let Assumption 3 on the data hold and consider the data-driven MPC scheme summarized in Algorithm 1 using the kEDMD surrogate model (21) of the control-affine dynamics (19) in the optimization step of Algorithm 1. Let $N\in\mathbb{N}$ be such that $\alpha_{N}$ defined in (11) is in $(0,1]$ and assume that the system is cost controllable on a set $S\subseteq\Omega\ominus\mathcal{B}_{r(N)\eta^{\varepsilon}}$, see Definition 2, with $\mathbf{u}\in\mathcal{U}_{N}$. Then, if the fill distance $h_{\mathcal{X}}$ and the cluster radius $r_{\mathcal{X}}$ are sufficiently small, Algorithm 1 asymptotically stabilizes the system (19).
Proof. We have verified (P-bound) and (U-bound) of Assumption 1, and Assumption 2 for the kEDMD surrogate of the control-affine dynamics (19) in Proposition 2, Theorem 3, and Lemma 1, respectively. Hence, all assumptions of Theorem 1 hold in view of the imposed cost controllability and the assertion can be inferred. $\square$
@@ 4.3. Relation to kernel approximation of control systems
$\widetilde{F}_{i}(x)=\sum_{j=1}^{d}(K_{\mathcal{X}}^{-1}(F_{i})_{\mathcal{X}})_{j}\Phi_{x_{j}}(x)=(F_{i})_{\mathcal{X}}^{\top}K_{\mathcal{X}}^{-1}\hskip-2.84526pt\begin{bmatrix}\Phi_{x_{1}}(x)\\ \vdots\\ \Phi_{x_{d}}(x)\end{bmatrix}$ (26)
for each $i\in[1:n]$. This corresponds to direct interpolation of the flow map and hence involves only one projection. Moreover, the main difference to the approximation (17) are the data requirements: For the approximation (17) applied to the coordinate functions, we require data through the lens of the features, i.e., $\mathsf{k}(x_{i},F(x_{j}))$ for $(i,j)\in[1:d]^{2}$, whereas for the approximation (26), we require samples of the flow map, i.e., $F(x_{j})$ for all $j\in[1:d]$. The error bound for the surrogate dynamics (26), however, is very closely related to the one inferred from (17) and of similar structure, see (Köhne et al., 2025, Theorem 3.4), where the first estimate in this reference corresponds to (26) (involving only one projection) and the second estimate to (17) (involving two projections). As both projection errors are proportional to the fill distance $h_{\mathcal{X}}$, the convergence rate is same and the results of Subsection 4.2 can be derived analogously. Hence, Assumption 1 can also be verified for the surrogate model (26).
@@ 5. Numerical simulations
In this section, we illustrate the asymptotic stability of the origin w.r.t. the MPC closed loop using the kernel-EDMD surrogate in the prediction and optimization step by conducting numerical simulations. In the following, we denote by physics-informed kEDMD (PI-kEDMD) the models obtained with the modified Step 1 for the encoding of the equilibrium as described in Section 4.2, and with kEDMD the models obtained according to Bold et al. (2025b) without modifying Step 1.
@@ 5.1. Van der Pol
with parameters $\Delta t=0.05$, $\nu=0.1$, and control constraints $\mathbb{U}=[-2,2]$ on the domain $\Omega=[-2,2]^{2}$. The kEDMD surrogate (21) is constructed using Wendland kernels with smoothness degree $k=1$. The cluster points are chosen according to a Padua grid. The set of Padua points of order $p$ in the domain $1,1]\times[-1,1]$ is defined as
@@
Moreover, we add a cluster point in the origin to the Padua grid, according to Assumption 3. The Padua grid is properly scaled to cover the domain $\Omega$. Padua grids of order $p\in\{25,50\}$ are considered, resulting in $d\in\{352,1327\}$ state-space sample points. In comparison to a uniform grid, the use of a Padua grid allows to have a more uniform error through the domain, and lower condition numbers in the model matrices. In fact, Padua points minimize, analogously to Chebychev nodes in one dimension, the Lebesgue constant and, thus, the condition number and are, thus, preferable. The use of a uniform grid with a large number of data points may lead to numerical problems in the solution of the optimal control problem.
For each cluster point $x_{i}\in\mathcal{X}$, $d_{i}=25$ random control values $u_{ij}\in\mathbb{U}$ are chosen, yielding the data points $(x_{ij},u_{ij},x_{ij}^{+})$. Herein, $x_{ij}$ and $u_{ij}$ are drawn according to our data requirements with $r_{\mathcal{X}}=\nicefrac{{\sqrt{2}}}{{d}}$ to specify the neighborhood of $x_{i}$, see Section 4. For the derivation of the PI-kEDMD model Step 1 of the algorithm is modified for the cluster point $x_{1}=0$ by setting $\tilde{g}_{0}(x_{1})=0$, while in the kEDMD model Step 1 is performed in the same way for all cluster points. The MPC cost function is chosen as in (4) with matrices $Q=I_{2}$ and $R=10^{-4}$. The following closed-loop simulations emanate from the initial state $x^{0}=(0.5,0.5)^{\top}$, and consider a prediction horizon $N\in\{10,30\}$ in MPC. The simulations are implemented in Python, and Casadi Andersson et al. (2019) is used for the solution of the MPC optimization problem. The comparison of the closed loop error with different number of clusters and different prediction horizons is reported in Fig. 1.
@@
In the simulations with the kEDMD model, we observe practical asymptotic stability only: the error stagnates at a positive constant value. The tracking accuracy increases with the number of cluster points. This is expected, as a large number of cluster points is related to lower modeling errors. On the other hand, the PI-kEDMD models allow to reach a lower error, with a decrease up to the solver tolerance at around $10^{-14}$, in accordance to the shown asymptotic stability property of the MPC closed loop. We also notice that a fast decrease to the origin is obtained with the prolonged prediction horizon $N=30$ – a property related to the degree of suboptimality $\alpha_{N}$. For both the values of the prediction horizon, the asymptotic behavior of the PI-kEDMD-based MPC is the same that is obtained in a nominal MPC using the real system equations in the optimal control problem.
@@
In Fig. 2, we report the trajectories of the same simulations in the phase space. We can see that, while the asymptotic behavior is strongly related to the accuracy of the model in the origin, the transient behavior is mainly influenced by the prediction horizon and the number of cluster points. Finally, Table 1 reports the average computation time for the solution of the MPC on a laptop with an Intel i7-1065G7 CPU. The computation times increase when considering larger numbers of cluster point and longer prediction horizons, and are comparable for the PI-kEDMD and the kEDMD model.
PI-kEDMD kEDMD
$N=10$ $N=30$ $N=10$ $N=30$
$d=352$ 0.039 0.114 0.049 0.118
$d=1327$ 0.146 0.397 0.141 0.403
Table 1. Average computation time in seconds of the kEDMD-based MPC for the Van der Pol oscillator, for different numbers of prediction horizons and of number of data points.
@@ 5.2. Four-tanks process
As a second example, we consider the four-tank system (see Alvarado et al. (2011)) described by the differential equation
@@
The state of the system is given by the levels in the four tanks, i.e. $x=(h_{1},\;h_{2},\;h_{3},\;h_{4})^{\top}\in\mathbb{R}^{4}$, while the control variables are the flows through the two valves, i.e. $u=(q_{a},\;q_{b})^{\top}\in\mathbb{R}^{2}$. The numerical values of the system parameters can be found in Alvarado et al. (2011) and as sampling time we choose $\Delta t=10s$. The discrete-time version of (28) obtained with the forward Euler method is considered as ground truth. The controlled state for this system is the equilibrium given by $\bar{x}=(0.65,0.66,0.6417,0.6882)^{\top}m$ and $\bar{u}=(1.666,1.974)^{\top}m^{3}/h$. To derive the kEDMD surrogate model, the state and inputs are shifted around the origin. The state domain, where we sample, is $\Omega=[0.2\ m,1.36\ m]^{2}\times[0.2\ m,1.30\ m]^{2}$, while the input is constrained by the set $\mathbb{U}=[0\ m^{3}/h,3.26\ m^{3}/h]\times[0\ m^{3}/h,4\ m^{3}/h]$. The cluster centers are chosen in a uniform grid in $\Omega$. Moreover, an additional cluster center is added at the equilibrium. We consider datasets with $d\in\{626,1297,2402\}$ cluster centers, that correspond to $\sqrt[4]{d-1}\in\{5,6,7\}$. We decided to stick to a uniform grid, because the explicit construction of a grid minimizing the Lebesgue constant in dimensions higher than two is more involved and we did not run into numerical issues. Possible approaches to reduce the condition number of the model matrices are the use of the cartesian product of Padua grids or of more sophisticated approaches as outlined, e.g., in Jiménez-Ramos et al. (2025). The cluster radius is $r_{\mathcal{X}}=2/d$ for all the models, and 25 data points $(x_{ij},u_{ij},x_{ij}^{+})$ are considered in each cluster. For the closed-loop simulations, we consider an initial state $x_{0}=(1.0,1.0,1.0,1.0)^{\top}m$, MPC cost matrices $Q=I_{4}$ and $R=10^{-4}I_{2}$, and a prediction horizon $N=10$.
@@
The errors of the closed loop simulations are reported in Fig. 3. As in the previous example, we can see that the error reaches a higher accuracy with the PI-kEDMD model, and the accuracy increases with a larger number of cluster points. The number of cluster points also influences how quickly the error decreases, which occurs more slowly for $d=626$. In this example, the closed loop behavior remains the same if the prediction horizon is increased to $N=20$. Finally, the average computation times for the solution of the MPC optimization problem are reported in Table 2.
PI-kEDMD KEDMD
$d=626$ 0.370 0.383
$d=1297$ 0.875 0.864
$d=2402$ 1.876 1.744
Table 2. Average computation time in seconds of the kEDMD-based MPC for the four-tank system, for different number of data points and prediction horizon $N=10$.
@@ 6. Conclusions and outlook
We have proposed a framework to rigorously show asymptotic stability of an equilibrium w.r.t. the MPC closed, where data-driven surrogate models are used in the optimization step assuming cost controllability of the original system. The key ingredients are three properties of the data-driven surrogate: First, Lipschitz continuity uniform in the approximation accuracy. Second and third, uniform and foremost proportional error bounds, see Assumption 1. Then, we rigorously verified all assumptions for a data-driven kEDMD surrogate models using Koopman operator theory resulting in finite-data error bounds. In particular, we have proposed a learning algorithm based on flexible data sampling. Finally, the theoretical results were verified by numerical simulations for two representative examples.
Future work could consider an extension towards Koopman-based control of partial differential equations Deutscher (2024) as well as an extension towards novel (E)DMD variants such as eigenfunction-based approximation via (control) Liouville operators Rosenfeld and Kamalapurkar (2024), residual DMD Colbrook et al. (2023) or via stochastic Koopman approximants Wanner and Mezić (2022) using recently proposed finite-data error bounds for kEDMD derived Hertel et al. (2025). Moreover, the stability analysis can be extended to consider the presence of state constraints, disturbances and time-varying reference signals, which are not taken into account in this paper and require additional treatment as outlined in Remark 1. In addition, also an extension towards input-output data based on, e.g., Manzano et al. (2020), would be of interest.